OpenStudy (owllover101):

Which ordered pairs are solutions to the equation? y = 2x + 5 A. (5, 15) and (2, 9) B. (5, 10) and (2, 9) C. (5, 15) and (9, 2) D. (15, 5) and (2, 9)

OpenStudy (owllover101):

ill give medal

OpenStudy (owllover101):

and will fan

OpenStudy (dtan5457):

Have you tried plugging in each point and seeing if the y values are true?

OpenStudy (owllover101):

i dont get this

OpenStudy (dtan5457):

Ok.. y =2x+5 To test the point (5, 15) 5=x y=2(5)+5 y=15, this point is a solution

OpenStudy (owllover101):

still confused

OpenStudy (owllover101):

???????????????

OpenStudy (dtan5457):

What don't you get..?

OpenStudy (owllover101):

basicly all of it

OpenStudy (dtan5457):

@Nnesha You have better explanation?

Nnesha (nnesha):

in order pair first number is x and 2nd one is y for example \[\huge\rm (x ,y)\] replace x and y by one of order pair if both sides are equal then that order pair iz ur answer you should try both.

OpenStudy (owllover101):

what do you mean if both sides are equal

OpenStudy (dtan5457):

your function y=2x+5...if both sides are equal..

OpenStudy (owllover101):

like.....

OpenStudy (owllover101):

i will never get this

Nnesha (nnesha):

both side equal mean 4 = 4 0 = 0 when you solve one of side that should be equal to the other side

OpenStudy (owllover101):

ohh

OpenStudy (owllover101):

so the answer would be ??????????????

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

Let's test the claim "(5,15) is a solution to the equation y = 2x+5" so that means (x,y) = (5,15) so x = 5 and y = 15